OpenStudy (anonymous):

The area of a triangle is 63 square inches. The base is 18 inches long. What is the height of the triangle?

DivineSolar (divinesolar):

Hold on please! :)

DivineSolar (divinesolar):

Dont countdown lol that wont get it done faster i wish yours was finding the base length. That would be easier.

OpenStudy (ilovewolf):

Are there options?

DivineSolar (divinesolar):

If it was i have my answer, Area of triangle = 1/2*base*height = 63 sqare inches. height = 18 inches Area = 1/2*base*18 = 63 base*18 = 63*2 base*18 = 126 base = 126/18 = 7 inches Answer: the length of base is 7 inches.

DivineSolar (divinesolar):

That would be the answer of the length base was the question.

DivineSolar (divinesolar):

Does that happen to be one of your questions?

OpenStudy (anonymous):

@ilovewolf if i put options, people would randomly put in letters or answers

DivineSolar (divinesolar):

Well put in options for me and him as i try to be as accurate as possible. Deny every elses answers if that's what you want to believe.

OpenStudy (anonymous):

there you go

OpenStudy (anonymous):

@DivineSolar @ilovewolf you still there?

OpenStudy (anonymous):

@DivineSolar please come back

DivineSolar (divinesolar):

I'm back. Firefox wouldn't work.

DivineSolar (divinesolar):

I don't really know how to explain it but is 7 one of your choices?

DivineSolar (divinesolar):

@jakematt24 You there?